OpenStudy (meowlover17):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8

OpenStudy (meowlover17):

−96 0 24 48

OpenStudy (meowlover17):

@090909090909

OpenStudy (anonymous):

Hi @MeowLover17 First, we need to find the value of g(6). We do this by substituting in 6 for x. g(6) = 6-8 = -2 Now that we know g(6) = (-2), we can substitute that into f(x). f(-2) = 6(-2) + 12 = -12 + 12 = 0. Let me know if there is anything you don't understand!